Do you now two numbers that are multipled to get -96, and added to get -46?
Answer by ankor@dixie-net.com(22740)

two numbers that are multiplied to get -96, and added to get -46?
Two numbers x & y, two equations
xy = -96
x+y= -46
y = (-46-x)
replace y in the 1st equation
x(-46-x) = -96
-46x - x^2 = -96
x^2 + 46x = 96; multiplied by -1
A quadratic equation
x^2 + 46x - 96 = 0
Factors to
(x+48)(x-2) = 0
two solutions
x = -48 then y = 2
and
x = +2 then y = -48